OpenStudy (anonymous):

determine the value of dy/dt at x=3 when y=x^2-5x and dx/dt=3

OpenStudy (anonymous):

Chain rule. \[\frac{dy}{dt}=\frac{dy}{dx} \frac{dx}{dt}\]

OpenStudy (anonymous):

ohhhh thank you!!